# A Weighted Loss Approach to Robust Federated Learning under Data Heterogeneity

### v2

@@ 3.2.3 Our proposition
While many approaches focus on handling update heterogeneity [3, 24, 4, 15, 12], WoLA tackles its root cause—label skewness—by aligning gradients from the start. This enables WoLA to significantly outperform state-of-the-art methods. In the next section, we explain the motivation behind its design.
4 Rationale for Worker Label Alignement Loss
This section first shows that local gradients approximate weighted averages of class-wise gradients. Next, it illustrates that class-wise gradients can differ significantly. Finally, it presents the main conclusion drawn from these observations.
4.1 Gradients as Weighted Averages
The key rationale behind WoLA is that the gradient over a sufficiently large local dataset can be approximated as a weighted average of class-wise gradients, with weights given by the local label distribution.
More precisely, the gradient of worker $i$ for a standard loss is:
$\nabla\mathcal{L}_{i}$ $=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
$=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\sum_{c=1}^{C}\mathbbm{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}$
$=\sum_{c=1}^{C}\frac{N^{c}_{i}}{N_{i}}\frac{\sum_{(x,y)\in\mathcal{D}_{i}}\mathbbm{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}}{N^{c}_{i}},$
where:
$\frac{N^{c}_{i}}{N_{i}}=\Bigg{(}\text{Proportion of class }c\text{ in }\mathcal{D}_{i}\Bigg{)}\vcentcolon=p^{c}_{i},$
and
$\frac{\sum_{(x,y)\in\mathcal{D}_{i}}\mathbbm{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}}{N^{c}_{i}}$ (1)
$\quad\quad=\Bigg{(}\text{Mean gradient of class }c\text{ in }\mathcal{D}_{i}\Bigg{)}\vcentcolon=\mu_{i}^{c}$
Given a sufficient number of instances of class $c\in[C]$, the following approximation is reasonable (see Appendix Section A.2, Lemma 2 for a formal statement):
(Mean gradient of class $\Bigg{(}\text{Mean gradient of class }c\text{ in }\mathcal{D}_{i}\Bigg{)}$ (2)
$\quad\quad\approx\Bigg{(}\text{Mean gradient of class }c\text{ in}\mathcal{D}\Bigg{)}$
$\quad\quad=\frac{\sum_{(x,y)\in\mathcal{D}}\mathbbm{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}}{N^{c}}\vcentcolon=\mu_{c}$
We call $\mu_{i}^{c}$ the class-$c$ gradient of worker $i$, and $\mu^{c}$ the global class-$c$ gradient. When clear from context, we omit the term global.
Using approximation equation (2), we can express the local gradient as a weighted average of global class gradients:
$\nabla\mathcal{L}_{i}\approx\sum_{c=1}^{C}p_{i}^{c}\mu^{c}$ (3)
Assuming that the vectors $\mu^{c}$ for $c\in[C]$ differ—either in norm or direction—the differences in local gradients arise from the local label distributions, through the weights $p_{i}^{c}$.
In the next subsection, we show that the vectors $\mu^{c}$ for $c\in[C]$ are indeed different.
4.2 Difference in Class Gradients
Figure 1 shows standard (non-FL) training of a shallow neural network on the Iris dataset [17]. The loss decreases and accuracy improves over epochs, indicating successful model convergence. Independently of the optimization, at each step we compute the class gradients for Setosa, Versicolour, and Virginica, and track the cosine similarity between these gradients over training steps.
The figure reveals that the angles between class-wise gradients often exceed $\pi/2$, indicating that the gradients point in widely different directions. In some cases, such as between Versicolour and Virginica, the gradients are nearly opposite.
This highlights that different classes can push the model in distinct—and sometimes conflicting—directions during training. Similar observations have been reported in previous studies [46, 4].
[…]
4.3 Key Takeaways
Local gradients approximate weighted averages of global class gradients, with weights given by the local label distribution. Since class gradients may differ, aligning local gradients requires aligning their weightings—as if local datasets shared the same label distribution. WoLA is designed to achieve this goal.
5 Worker Label Alignement Loss
This section introduces WoLA. It first defines the training objective required for WoLA, then presents its formula. The final subsections provide an intuitive understanding of WoLA’s properties. These properties are rigorously stated in Section 6 and proved in the Appendix A.
5.1 Shared Training Objective
WoLA requires a shared training objective across workers. This objective is formalised by a probability distribution $q=(q^{1},q^{2},\ldots,q^{C})$, where $q^{i}\in[0,1]$ and $\sum_{i\in[C]}q^{i}=1$. A larger $q^{i}$ means that point-wise errors for samples with label $i$ are given more weight during training, and vice versa. To define this distribution, we considered three situations:
@@ 5.1 Shared Training Objective
Local label distributions are shared
Each worker shares its local label distribution with the server before training. The server then computes the global distribution. In this case, it is important to note that (i) the process reduces data privacy, and (ii) it introduces a new opportunity for Byzantine workers to attack the system. We carefully evaluate this attack scenario in Section 7.2. Regarding privacy concerns, we assume in this paper that the server is trusted as we consider privacy as an orthogonal problem that is actively being investigated in the literature. Nevertheless, off-the-shelf SecureAggregation protocols such as [6] could be easily used to release the trusted server assumption.
@@
All workers agree that the global model should favor certain classes during training, regardless of the actual global label distribution. For example, in a task to detect hateful or illegal content on social networks—where hate content is rare—it may be agreed to assign higher importance to these cases. In medical diagnosis, rare diseases may also be given more weight. In contrast, for automated recruitment systems, all classes may be weighted equally to avoid algorithmic bias.
The training objective $q$ is shared by the server with all workers once before training begins, allowing them to adjust their local loss functions accordingly.
5.2 WoLA definition
@@
WoLA locally adjusts the weights of point-wise errors to simulate an identical label distribution across workers. More precisely, WoLA for worker $i$, denoted $\mathcal{WL}_{i}$, is defined as:
$\mathcal{WL}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\frac{q^{y}}{p^{y}_{i}}\ell\Big{(}y,\Phi(x)\Big{)},$
While the formula is undefined for $p^{y}_{i}=0$, it is not problematic, as it indicates that label $y$ is absent from the local dataset $\mathcal{D}_{i}$, meaning the corresponding weight never appears in the sum.
A sample from class $c$ is overweighted when class $c$ is underrepresented locally compared to the training objective, i.e., when $p_{i}^{c}<q^{c}\Leftrightarrow 1<q^{c}/p_{i}^{c}$; and underweighted otherwise.
5.3 Local Updates Alignement
WoLA makes local datasets appear as if they had identical label distributions matching the training objective $q$—that is, as if $p_{i}^{c}=q^{c}$ for all $c\in[C]$ and all $i\in\mathcal{H}$. Specifically, for worker $i$, the gradient under WoLA is:
$\nabla\mathcal{WL}_{i}\vcentcolon=$ $\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\frac{q^{y}}{p^{y}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
$=$ $\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\sum_{c=1}^{C}\mathbbm{1}(y=c)\frac{q^{c}}{p^{c}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
$=$ $\sum_{c=1}^{C}q^{c}\frac{\sum_{(x,y)\in\mathcal{D}_{i}}\mathbbm{1}(y=c)\nabla\ell\Big{(}y,\Phi(x)\Big{)}}{N^{c}_{i}},$
using approximation in equation (2), it follows:
$\nabla\mathcal{WL}_{i}\approx\sum_{c=1}^{C}q^{c}\mu^{c}$ (4)
With WoLA, the approximation of the worker $i$’s gradient no longer depends on local label proportions. It becomes a weighted average of class gradients, using the shared training objective $q$ as weights. Since equation (4) holds for all $i\in[C]$, it follows that $\nabla\mathcal{WL}_{i}\approx\nabla\mathcal{WL}_{j}$ for all $j\in[C]$. This suggests that local updates are more similar under WoLA, potentially reducing gradient dissimilarity and improving robustness.
Figure 2 illustrates the impact of WoLA on updates using a fictitious example with three classes: Dog, Flower, and Frog, under the training objective $q=(1/3,1/3,1/3)$. It shows three local datasets with label distributions $(3/6,2/6,1/6)$, $(2/6,2/6,2/6)$, and $(1/6,1/6,4/6)$. Two of them differ from the training objective $q$. WoLA addresses this mismatch by upweighting samples from local minority classes and downweighting those from local majority classes, thereby aligning updates.
(a) Point-wise standard loss gradients on $\mathcal{D}_{1}$
(b) Point-wise standard loss gradients on $\mathcal{D}_{2}$
(c) Point-wise standard loss gradients on $\mathcal{D}_{3}$
(d) Point-wise WoLA gradients on $\mathcal{D}_{1}$
(e) Point-wise WoLA gradients on $\mathcal{D}_{2}$
(f) Point-wise WoLA gradients on $\mathcal{D}_{3}$
Figure 2: Fictitious FL classification task with three classes (Dog, Flower, Frog) under the training objective $q=(1/3,1/3,1/3)$, comparing standard loss with WoLA. Each image represents a data point from a local dataset. Arrows indicate 2D point-wise loss gradients; in black squares, arrows show the resulting updates. Subfigures (a–c) display gradients under the standard loss, while (d–f) show gradients with WoLA. In (d) and (f), dataset sizes are scaled to reflect the effect of WoLA weights, although datasets are actually unchanged. For example, $\mathcal{D}_{3}$ has a local label distribution of $(1/6,1/6,4/6)$, while $q$ is uniform. This results in weights: $1/3\div 1/6=4/2$ for Dog and Flower, and $1/3\div 4/6=1/2$ for Frog. To avoid fractional images, the denominator 2 is omitted, giving 4 duplicates of Dog and Flower gradients, and 1 of Frog. The same logic applies to $\mathcal{D}_{1}$. Subfigures (b) and (e) are identical, as $\mathcal{D}_{2}$ label distribution matches the training objective. Local updates with WoLA are more aligned than with the standard loss.
5.4 Targeted Model
Assuming that the training objective $q$ matches the global label distribution $p$, i.e., $p=q$, the average gradients of the standard loss and the WoLA loss—denoted $\nabla\mathcal{L}$ and $\nabla\mathcal{WL}$ respectively—are similar. Using the approximations from equations (3) and (4), we have:
$\nabla\mathcal{L}$ $=\sum_{i\in\mathcal{H}}\nabla\mathcal{L}i\frac{N_{i}}{N}$
$\approx\sum_{i\in\mathcal{H}}\sum_{c=1}^{C}p_{i}^{c}\mu^{c}\frac{N_{i}}{N}=\sum_{c=1}^{C}p^{c}\mu^{c}=\sum_{i\in\mathcal{H}}\sum_{c=1}^{C}p^{c}\mu^{c}\frac{N_{i}}{N}$
$=\sum_{i\in\mathcal{H}}\sum_{c=1}^{C}q^{c}\mu^{c}\frac{N_{i}}{N}\approx\sum_{i\in\mathcal{H}}\nabla\mathcal{WL}_{i}\frac{N_{i}}{N}=\nabla\mathcal{WL}$
Therefore, when $p=q$, the training targets the same model regardless of whether the standard or WoLA loss is used.
5.5 Key Takeaways
WoLA requires a shared training objective. Previous analyses suggest that: (i) using WoLA reduces gradient dissimilarity, and (ii) when the training objective matches the label distribution, the model targeted with WoLA is the same as the one with the standard loss.
6 Theoretical Arguments
This section uses probabilistic tools to rigorously state the properties of WoLA. We first describe the data modeling framework, then present asymptotic results that offer insights into WoLA’s behavior.
6.1 Data Modeling
@@
Let $(X_{1},Y_{1},Z_{1}),(X_{2},Y_{2},Z_{2}),\ldots,(X_{N},Y_{N},Z_{N})$ be independent and identically distributed (i.i.d.) random triplets drawn from a distribution $\mathfrak{D}$ over the sample space $\mathcal{X}\times[C]\times\mathcal{H}$. Each pair $(X_{i},Y_{i})$ represents a datapoint in the dataset $\mathcal{D}$. For each $k\in[N]$, the random variable $Z_{k}$ takes values in $\mathcal{H}$ and indicates the worker to which the $k$-th datapoint is assigned. We also denote by $(X,Y,Z)$ a generic triplet with the same distribution $\mathfrak{D}$.
@@
For any triplet $(X_{k},Y_{k},Z_{k})$, we assume that $X_{k}$ and $Z_{k}$ are conditionally independent given $Y_{k}$. That is, conditioned on the label, the features provide no information about the assignment, and vice versa. Formally, $P(X_{k},Z_{k}|Y_{k})=P(X_{k}|Y_{k})P(Z_{k}|Y_{k}).$
6.2 Theoretical Results
The results are asymptotic, considering the case where the size of the global dataset tends to infinity, i.e., $N\to\infty$, and are based on the Law of Large Numbers. Full proofs and details are given in Appendix A.
@@ 6.2 Theoretical Results
Let $L_{i}$ and $W_{i}$ be the random counterparts of $\nabla\mathcal{L}_{i}$ and $\nabla\mathcal{WL}_{i}$, respectively. Let $L$ and $W$ denote the average gradients of honest workers for the standard loss and WoLA, respectively.
@@
With the standard loss, local gradients remain different even with an infinite amount of data:
@@
[…]
@@
[…]
In contrast, with WoLA, local gradients become asymptotically identical, regardless of the training objective shared by the server:
@@
[…]
@@
[…]
This directly implies that, under WoLA, the gradient dissimilarity vanishes asymptotically as $N\to\infty$ (see Appendix A.5).
Finally, if the training objective matches the global label distribution, i.e., $q=p$, then the standard loss gradient and the WoLA gradient become asymptotically identical:
@@
[…]
This equality implies that, in the asymptotic regime, when the training objective matches the label distribution, the model targeted with WoLA is the same as the one with the standard loss.
(a) Results under ALIE attack.
(b) Results under LF attack.
[…]
(a) Boxplot of accuracy gains based on Table 2.
(b) Boxplot of accuracy gains under the strongest attack, based on Table 3.
[…]
Table 2: Global Performance: Test Accuracy (s.d.) %, averaged over training. Results are averaged across aggregators (CWMed, CwTM, GM, MKrum) and attacks (ALIE, FOE, LF, Mimic, SF). Colors rank values within columns: yellow (1st), blue (2nd), red (3rd). See full table in Appendix.
(a) MNIST
Accuracy
Accuracy Gain Over Top Non-WoLA Methods
WoLA 0.0 0.0 0.0 +6.4 0.0 +0.4 +2.4 +17.3 -0.4 +1.5 +21.4 +16.5
WoLA+NNM 0.0 +0.2 +0.6 +8.8 +0.1 +0.6 +3.1 +19.9 -0.2 +1.8 +22.5 +26.6
WoLA 0.0 0.0 -0.2 +5.9 -0.1 +0.3 +2.0 +13.6 -0.4 +1.4 +20.5 +9.3
WoLA+NNM 0.0 +0.1 +0.3 +8.3 -0.1 +0.5 +2.8 +10.3 -0.2 +1.8 +22.3 +10.7
(b) Fashion MNIST
@@
(c) Purchase100
Accuracy
Accuracy Gain Over Top Non-WoLA Methods
WoLA -1.0 +1.3 +5.4 +14.6 -1.4 +3.1 +9.3 +14.8 -6.4 -1.7 +14.6 +6.8
WoLA -1.2 +0.8 +4.0 +10.2 -1.6 +2.4 +7.2 +5.2 -6.7 -2.9 +11.9 -3.9
(d) CIFAR10
@@
Table 3: Worst-case Performance: Test Accuracy (s.d.) % under the worst attack, averaged over training. Results are averaged across aggregators (CWMed, CwTM, GM, MKrum). Colors rank values within columns: yellow (1st), blue (2nd), red (3rd). See full table in Appendix.
(a) MNIST
Accuracy
Accuracy Gain Over Top Non-WoLA Methods
WoLA -0.1 +0.1 -0.8 +19.7 -0.1 +1.3 +6.0 +41.9 -0.4 +6.5 +28.1 +10.7
WoLA+NNM 0.0 +0.3 +1.3 +23.5 0.0 +1.6 +8.3 +44.3 -0.1 +7.0 +29.3 +6.0
WoLA -0.1 0.0 -0.8 +20.3 -0.1 +1.3 +5.7 +41.3 -0.4 +6.4 +26.3 +6.5
WoLA+NNM +0.1 +0.2 +0.9 +23.2 -0.1 +1.5 +7.9 +28.9 -0.1 +7.0 +30.6 -3.2
(b) Fashion MNIST
@@
(c) Purchase100
Accuracy
Accuracy Gain Over Top Non-WoLA Methods
WoLA -0.7 +3.4 +9.2 +29.0 +0.8 +10.0 +15.4 +16.1 -1.5 +10.4 +15.1 +3.6
WoLA -0.9 +2.8 +7.6 +23.2 +0.5 +9.1 +12.1 +18.4 -1.8 +8.9 +20.3 +1.7
(d) CIFAR10

### v3

@@ Introduction
While many approaches address update heterogeneity (Allouah et al. 2023; Karimireddy, He, and Jaggi 2020; Bao, Wu, and He 2024; Fang et al. 2025; Data and Diggavi 2021), WoLA tackles its root cause—label skewness—by aligning gradients from the start.
Worker Label Alignement Loss
This section introduces WoLA. We first present the rationale behind its design, then provide its formal definition and training objective. The final subsections describe the theoretical framework and the properties of WoLA. Proofs and additional details are provided in the Appendix.
Rationale for WoLA
We begin by observing that local gradients are weighted averages of class-wise gradients:
$\nabla\mathcal{L}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}=\sum_{c=1}^{C}p_{i}^{c}\mu^{c}_{i},$
where $p^{c}_{i}\vcentcolon=N^{c}_{i}/N_{i}$ is the proportion of class $c$ in $\mathcal{D}_{i}$, and
$\mu_{i}^{c}\vcentcolon=\frac{\sum_{(x,y)\in\mathcal{D}_{i}}\mathbf{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}}{N^{c}_{i}}$
is the average gradient for class $c$ in worker $i$’s local dataset.
With sufficient samples per class, local class-wise gradients $\mu_{i}^{c}$ approximate the global class-wise gradients $\mu^{c}$:
$\nabla\mathcal{L}_{i}=\sum_{c=1}^{C}p_{i}^{c}\mu^{c}_{i}\approx\sum_{c=1}^{C}p_{i}^{c}\mu^{c},$ (1)
where the global class-wise gradient is defined as:
$\mu_{c}\vcentcolon=\frac{\sum_{(x,y)\in\mathcal{D}}\mathbf{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}}{N^{c}}$
In general, global class-wise gradients are not equal (Zhang et al. 2022; Bao, Wu, and He 2024); we illustrate this in the Appendix.
Assuming the above approximation in (1) holds, differences in local gradients primarily stem from differences in local label distributions (via $p_{i}^{c}$). Hence, aligning local gradients requires aligning local label distributions or simulating such alignment. WoLA is specifically designed to achieve this goal.
Training Objective & WoLA Definition
WoLA requires a training objective $q=(q^{1},q^{2},\ldots,q^{C})$ with $\sum_{i\in[C]}q^{i}=1$ and $q^{i}\geq 0$, which is shared with all workers once before training begins. This vector $q$ represents the simulated local label distribution of each worker when using WoLA. In practice, a higher $q^{i}$ increases the weight given to errors on class $i$ during local training.
The distribution $q$ may approximate the true global label distribution or be set arbitrarily. We present three settings for defining $q$:
Local label distributions are shared
Each worker communicates its local label distribution to the server, which computes and broadcasts the global label distribution. However, this introduces a potential vulnerability: Byzantine workers can manipulate the training objective by reporting false distributions. We analyze this attack in the Appendix (see Section Training Objective Attack).
@@
All workers may agree on a predefined target distribution, regardless of the true label distribution. For example, in fair hiring systems, classes may be weighted equally to mitigate algorithmic bias.
@@
WoLA adjusts point-wise error contributions to simulate a common label distribution $q$ across all workers.
Definition 1
The local WoLA loss for worker $i$, denoted $\mathcal{W}\!\mathcal{L}_{i}$, is defined as:
$\mathcal{W}\!\mathcal{L}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\frac{q^{y}}{p^{y}_{i}}\ell\Big{(}y,\Phi(x)\Big{)},$
A sample from class $c$ is up-weighted when class $c$ is underrepresented in the local dataset compared to the training objective $q$, and down-weighted otherwise. This formula is directly derived from Importance Sampling (Tokdar and Kass 2010); further details are provided in the Appendix.
Before establishing the theoretical properties that explain WoLA’s efficiency, we introduce the theoretical framework.
Theoretical Framework
@@
Let $(X_{1},Y_{1},Z_{1}),(X_{2},Y_{2},Z_{2}),\ldots,(X_{N},Y_{N},Z_{N})$ be independent and identically distributed (i.i.d.) samples from a distribution $\mathfrak{D}$ over $\mathcal{X}\times[C]\times\mathcal{H}$. Each pair $(X_{i},Y_{i})$ represents a datapoint in $\mathcal{D}$. For each $k\in[N]$, the random variable $Z_{k}$ takes values in $\mathcal{H}$ and indicates the worker to which the $k$-th datapoint is assigned. Let $(X,Y,Z)$ be a generic triplet with the same distribution $\mathfrak{D}$.
@@
For any triplet $(X_{k},Y_{k},Z_{k})$, we assume $X_{k}$ and $Z_{k}$ are conditionally independent given $Y_{k}$, i.e., $P(X_{k},Z_{k}|Y_{k})=P(X_{k}|Y_{k})P(Z_{k}|Y_{k})$. Given the label, features provide no information about assignments, and vice versa. This implies $P(X_{k}|Y_{k},Z_{k})=P(X_{l}|Y_{l},Z_{l})$ for any $k,l$, so the feature distribution is identical across workers.
WoLA Properties
@@
Let $L_{i}$ and $W_{i}$ be the random counterparts of $\nabla\mathcal{L}_{i}$ and $\nabla\mathcal{W}\!\mathcal{L}_{i}$, respectively. Let $L$ and $W$ denote the average gradients of honest workers for the standard loss and WoLA, respectively.
@@
Without WoLA, local gradients remain different even with an infinite amount of data.
@@
$L_{i}\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Z=i\Big{]}$
@@
$L\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{]}$
Under label skewness, and since class-wise gradients differ, there exists at least one worker $i$ for which the expected gradient differs from the global expectation:
$\mathbb{E}[\nabla\ell(Y,\Phi(X))|Z=i]\neq\mathbb{E}[\nabla\ell(Y,\Phi(X))]$
This implies $\lim L_{i}\overset{\text{a.s.}}{\neq}\lim L$ as $N\to\infty$, so the gradient dissimilarity remains strictly positive.
With WoLA, local gradients become asymptotically identical, regardless of the training objective shared by the server.
@@
$W_{i}\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}\sum_{c=1}^{C}q^{c}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
@@
$W\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}\sum_{c=1}^{C}q^{c}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
As a result, the gradient dissimilarity under WoLA vanishes as $N\to\infty$.
When the training objective matches the global label distribution (i.e., $q=p$), the WoLA and non-WoLA gradients become asymptotically identical.
@@
$\lim_{N\to\infty}W\overset{\text{a.s.}}{=}\lim_{N\to\infty}L$
In the asymptotic regime, when the training objective matches the global label distribution, WoLA and non-WoLA gradients target the same model.
Table 1: Test accuracy (± standard deviation) in %, averaged over training, on Fashion MNIST and CIFAR10. Global results are averaged across aggregators (CWMed, CwTM, GM, MKrum) and attacks (ALIE, FOE, LF, Mimic, SF). Worst-case results correspond to the worst attack, averaged across aggregators. Colors rank values within columns: yellow (1st), blue (2nd), red (3rd). See full table in Appendix.
(a) Fashion MNIST (Global)
@@
(b) CIFAR10 (Global)
@@
(c) Fashion MNIST (Worst-case)
@@
(d) CIFAR10 (Worst-case)